# Past Movements-Guided Motion Representation Learning for Human Motion Prediction

### v1

@@ 5 Ablation Study
In this section, we explore the impact of components of our network, including the mask strategy, pretraining paradigm. We provide studies on model architecture in the supplementary materials. All experiments were carried out on the Human3.6M dataset.
@@
Tab. 6 presents the results of two cases: A, without utilizing the velocity-based mask strategy (i.e. randomly mask 75% of the joints), and F, which represents the complete model. It can be seen that using the velocity-based mask reduced the average prediction errors by 2.3%. Tab. 6 also shows the comparisons of different mask rates (ranging from 0% to 75%), represented by B, C, D, F. The high mask rate facilitates the model to learn the space-time representation during the self-reconstruction process.
@@
We show the results of pretraining the model by denoising in Tab. 6 E. Compared with the self-reconstruction, denoising pretraining method lacks the exploration of the spatiotemporal relationship between joints.
Table 6: Impact of mask strategy and pretraining paradigm. $r$ is the mask rate. ’denoise’ means we conduct self-supervised pretraining by denoising.
@@ 6 Conclusion
In this paper, we propose a novel motion prediction paradigm, which consists two stages: 1) Past motion guided future motion self-reconstruction for learning the spatio-temporal representation of human joints (i.e. pretrain stage). 2) Train the network for the motion prediction task (i.e. finetune stage). Despite the simplicity of the model structure, our approach achieves the best results on three representative datasets and is highly extensible.
@@
Motion Walking Eating Smoking Discussion Direction
DMGNN[20] 17.3 30.7 54.6 65.2 11.0 21.4 36.2 43.9 9.0 17.6 32.1 40.3 17.3 34.8 61.0 69.8 13.1 24.6 64.7 81.9
MSR[7] 12.2 22.7 38.6 45.2 8.4 17.1 33.0 40.4 8.0 16.3 31.3 38.2 12.0 26.8 57.1 69.7 8.6 19.7 43.3 53.8
Mixer[2] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1 7.5 18.4 43.9 55.3
PGBIG[26] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7 7.2 17.6 40.9 51.5
SPGSN[24] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1 7.4 17.2 40.0 50.3
siMLPe[12] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1 11.0 19.4 41.6 52.1
DSTD-GC[10] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4 6.9 17.4 40.1 51.7
AuxFormer[41] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2 6.8 17.0 40.3 51.6
CIST-GCN[33] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3 7.3 18.1 43.6 55.3
Motion Greeting Phoning Posing Purchases Sitting
DMGNN[20] 23.3 50.3 107.3 132.1 12.5 25.8 48.1 58.3 15.3 29.3 71.5 96.7 21.4 38.7 75.7 92.7 11.9 25.1 44.6 50.2
MSR[7] 16.5 37.0 77.3 93.4 10.1 20.7 41.5 51.3 12.8 29.4 67.0 85.0 14.8 32.4 66.1 79.6 10.5 22.0 46.3 57.8
Mixer[2] 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5 13.4 30.7 65.6 80.0 9.8 20.9 45.6 57.3
PGBIG[26] 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6 12.5 28.7 60.1 73.3 8.8 19.2 42.4 53.8
SPGSN[24] 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5 12.7 28.6 61.0 74.4 9.3 19.4 42.2 53.6
siMLPe[12] 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6 15.5 29.3 59.3 72.0 13.2 21.9 44.5 55.7
DSTD-GC[10] 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3 12.7 29.7 62.3 75.8 8.8 19.3 42.9 54.3
AuxFormer[41] 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0 11.9 28.0 61.8 76.3 8.7 19.0 42.1 53.3
CIST-GCN[33] 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0 13.3 30.2 63.0 77.3 8.9 19.4 42.3 53.6
DMGNN[20] 15.0 32.9 77.1 90.3 13.6 29.0 46.0 58.8 12.2 24.2 59.6 77.5 47.1 93.3 160.1 171.2 14.3 26.7 50.1 63.2
MSR[7] 16.1 31.6 62.5 76.8 9.9 21.0 44.6 56.3 10.7 23.1 48.3 59.2 20.7 42.9 80.4 93.3 10.6 20.9 37.4 43.9
Mixer[2] 15.3 30.5 62.7 78.0 9.0 19.9 44.7 56.9 9.9 22.0 49.7 62.5 21.1 43.9 84.7 101.2 9.9 20.4 39.3 47.3
PGBIG[26] 13.9 27.9 57.4 71.5 8.4 18.9 42.0 53.3 8.9 20.1 43.6 54.3 18.8 39.3 73.7 86.4 8.7 18.6 34.4 41.0
SPGSN[24] 14.2 27.7 56.7 70.7 8.8 18.9 41.5 52.7 9.2 19.8 43.1 54.1 - - - - 8.9 18.2 33.8 40.9
siMLPe[12] 18.0 29.8 57.7 71.2 13.1 21.3 42.9 53.6 11.9 21.5 44.4 55.0 21.8 40.6 74.6 87.0 11.9 19.8 34.2 39.8
DSTD-GC[10] 14.1 28.0 57.3 71.2 8.4 18.8 42.0 53.5 8.7 20.2 44.3 55.3 19.6 41.8 77.6 90.2 9.1 19.8 36.3 42.7
AuxFormer[41] 13.5 27.6 57.7 72.2 8.2 18.4 41.5 53.0 8.2 18.5 41.2 52.2 17.1 36.5 70.4 83.0 7.8 15.9 30.2 37.0
CIST-GCN[33] 14.1 29.8 57.3 69.8 8.2 18.4 40.6 51.8 8.6 19.4 43.5 54.8 20.0 41.4 73.7 85.1 9.6 20.3 38.2 45.6
In this supplementary material, we provide more information that could not be included in the main manuscript because of space limit, including explanation of ST-Attention and PMG-Attention (in Sec. 7), implementation details (in Sec. 8), complete results for testing on Human3.6M dataset (in Sec. 9), ablation study on model architecture (in Sec. 10), and visualization of feature map (in Sec. 11).
7 Detailed Explanation of ST-Attention and PMG-Attention
In our model architecture for human motion prediction, both the Past Motion Encoder (PME) and the Future Motion Predictor (FMP) modules rely on attention mechanisms: SpatioTemporal Attention (ST-Attention) and Past Movement Guided Attention (PMG-Attention). These mechanisms are pivotal for capturing and integrating spatiotemporal dependencies in motion sequences.
ST-Attention
ST-Attention combines temporal and spatial attention mechanisms to model joint relationships over time effectively. This attention mechanism is utilized within PME to encode past motion features.
The temporal attention mechanism operates on the sequence dependencies over time in the joint positions:
$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{l}\right)\$ (10)
@@ 7 Detailed Explanation of ST-Attention and PMG-Attention
$\mathbf{Z}^{t}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{t}\left(\mathbf{K}^{t}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}$ (11)
where $\mathbf{H}^{l}$ represents the input at the $l^{th}$ layer, $\mathbf{Q}^{t},\mathbf{K}^{t}$, and $\mathbf{V}^{t}$ are queries, keys, and values, respectively, derived from $\mathbf{H}^{l}$, and $d_{k}$ is the dimension of keys.
The spatial attention mechanism focuses on modeling joint relationships in space:
$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}\left(\mathbf{Z}^{t}\right)$ (12)
@@
$\mathbf{Z}^{s}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{s}\left(\mathbf{K}^{s}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s}$ (13)
where $\mathbf{Q}^{s},\mathbf{K}^{s}$, and $\mathbf{V}^{s}$ are queries, keys, and values for spatial attention, respectively, similarly derived from $\mathbf{Z}^{t}$. $\mathbf{Z}^{s}$ is the output of ST-Attention.
PMG-Attention
PMG-Attention is employed within the Future Motion Predictor (FMP) module to incorporate past motion features into the prediction of future motion sequences. Unlike ST-Attention which focuses on the past motion sequence itself, PMG-Attention applies cross-attention between the current state and the past motion features:
$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{l}\right),\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{P}\right)$ (14)
@@
$\mathbf{Z}^{t}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{t}\left(\mathbf{K}^{t}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}$ (15)
$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}\left(\mathbf{Z}^{t}\right),\mathcal{F}^{s}_{QKV}\left(\mathbf{H}^{P}\right)$ (16)
@@
$\mathbf{Z}^{s}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{s}\left(\mathbf{K}^{s}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s}$ (17)
where $\mathbf{Q}^{t}$ and $\mathbf{Q}^{s}$ are obtained from the hidden state of future motions, and the $\mathbf{K}^{t},\mathbf{V}^{t},\mathbf{K}^{s},\mathbf{V}^{s}$ are obtained from the features of past motion.
@@ 8 Implementation Details
We deployed our network using an NVIDIA RTX4090 GPU (with PyTorch 2.0). For training, we set the batch size to 24 and utilized the Adam optimizer with an initial learning rate of 0.0005. To enhance convergence, we implemented a cosine annealing learning rate schedule. These training parameters were consistently applied across all datasets in our experiments. We ensured that the GPU memory usage remained below 6GB, enabling our network to be seamlessly deployed on various hardware platforms with ease.
9 Complete Results on Human3.6M

### v2

@@ 5 Ablation Study
We explore the impact of components of our network, including the mask strategy, pretraining paradigm. We provide studies on model architecture in the supplementary materials. All experiments were carried out on the Human3.6M dataset.
@@
Tab. 8 presents the results of two cases: A, without utilizing the velocity-based mask strategy (i.e. randomly mask 75% of the joints), and F, which represents the complete model. It can be seen that using the velocity-based mask reduced the average prediction errors by 2.3%. Tab. 8 also shows the comparisons of different mask rates (ranging from 25% to 75%), represented by B, C, F. The high mask rate facilitates the model to learn the space-time representation during the self-reconstruction process.
@@
We show the results of pretraining the model by denoising in Tab. 8 E. Compared with the self-reconstruction, denoising pretraining method lacks the exploration of the spatiotemporal relationship between joints. The result without pretraining is denoted as D, which clearly demonstrates the crucial role of our pretraining strategy in improving prediction performance.
Table 8: Impact of mask strategy and pretraining paradigm. $r$ is the mask rate. “denoise” means we conduct self-supervised pretraining by denoising.
@@ 6 Conclusion
In this work, we addressed the core limitations of end-to-end regression in motion prediction by introducing a two-stage framework that separates representation learning from prediction. Through past-guided future reconstruction and velocity-based masking, our self-supervised pretraining effectively captures intrinsic spatiotemporal dependencies. The joint fine-tuning of motion prediction and semantic future understanding further enhances coherence and realism.
[…]
@@
Figure 7: Visualization of motion prediction results on the processed FineMotion dataset.
Figure 8: Visualization of motion prediction results (“Posing”, “Walking”, “Greeting”) on Human3.6M dataset.
Figure 9: Visualization of motion prediction results (“Posing”, “Walking”, “Greeting”) on Human3.6M dataset.
In this supplementary material, we provide additional technical details, extended results, and further qualitative analyses that complement the main paper. Specifically, we include a detailed explanation of the spatiotemporal attention mechanisms used in our architecture (Sec. 7), implementation details and training configurations (Sec. 8), full motion prediction results on the Human3.6M dataset (Sec. 13), comprehensive ablation studies on model components and architectural choices (Sec. 10), and extensive visualizations, including motion prediction results, and attention feature maps (Sec. 9 and Sec. 12). We also discuss the mask strategy and reconstruction performance in Sec. 11. Together, these supplementary materials offer a deeper understanding of our framework and provide further evidence of its effectiveness and robustness.
7 Detailed Explanation of Model Architecture
We provide additional details of the two attention modules used in our framework: the SpatioTemporal Attention (ST-Attention) in the Past Motion Encoder (PME), and the spatiotemporal cross-attention in the Future Motion Decoder (FMD). Both modules follow the standard Transformer design, where each attention sub-layer (temporal or spatial) is immediately followed by a position-wise feed-forward network (FFN), each equipped with its own residual connection and layer normalization.
ST-Attention (within PME).
Each ST-Attention layer contains a temporal attention block and a spatial attention block, applied sequentially. Given the input to the $l^{th}$ layer $\mathbf{H}^{l}$, the temporal self-attention is computed as:
$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}(\mathbf{H}^{l}),$ (11)
@@ 7 Detailed Explanation of Model Architecture
$\mathbf{A}^{t}=\operatorname{softmax}\!\left(\frac{\mathbf{Q}^{t}(\mathbf{K}^{t})^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}.$ (12)
A position-wise FFN follows:
$\mathbf{H}^{l,t}=\mathrm{LN}\!\left(\mathbf{A}^{t}+\mathrm{FFN}(\mathbf{A}^{t})\right).$ (13)
The refined features $\mathbf{H}^{l,t}$ then pass through a spatial attention block:
$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}(\mathbf{H}^{l,t}),$ (14)
@@
$\mathbf{A}^{s}=\operatorname{softmax}\!\left(\frac{\mathbf{Q}^{s}(\mathbf{K}^{s})^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s},$ (15)
followed again by an FFN:
$\mathbf{H}^{l+1}=\mathrm{LN}\!\left(\mathbf{A}^{s}+\mathrm{FFN}(\mathbf{A}^{s})\right).$ (16)
Spatiotemporal Cross-Attention (within FMD).
FMD incorporates the encoded past motion features $\mathbf{H}^{P}$ via spatiotemporal cross-attention, where queries come from the current decoding state and keys/values come from the past sequence. As in PME, each temporal and spatial block has an independent FFN.
Temporal cross-attention is given by:
$\mathbf{Q}^{t}=\mathcal{F}^{t}_{Q}(\mathbf{H}^{l}),\qquad\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{KV}(\mathbf{H}^{P}),$ (17)
@@
$\mathbf{A}^{t}=\operatorname{softmax}\!\left(\frac{\mathbf{Q}^{t}(\mathbf{K}^{t})^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t},$ (18)
followed by:
$\mathbf{H}^{l,t}=\mathrm{LN}\!\left(\mathbf{A}^{t}+\mathrm{FFN}(\mathbf{A}^{t})\right).$ (19)
Spatial cross-attention further injects structural cues from past motion:
$\mathbf{Q}^{s}=\mathcal{F}^{s}_{Q}(\mathbf{H}^{l,t}),\qquad\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{KV}(\mathbf{H}^{P}),$ (20)
@@
$\mathbf{A}^{s}=\operatorname{softmax}\!\left(\frac{\mathbf{Q}^{s}(\mathbf{K}^{s})^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s},$ (21)
with its FFN:
$\mathbf{H}^{l+1}=\mathrm{LN}\!\left(\mathbf{A}^{s}+\mathrm{FFN}(\mathbf{A}^{s})\right).$ (22)
This spatiotemporal cross-attention design enables FMD to effectively leverage both temporal evolution and spatial structure in the past sequence when generating future motions.
@@ 8 Implementation Details
We implement all models using PyTorch 2.0 and train them on a single NVIDIA RTX4090 GPU. The batch size is set to 24, and the Adam optimizer is used with an initial learning rate of 0.0005. A cosine annealing schedule is adopted to gradually decay the learning rate, and the overall GPU memory consumption remains below 6GB, ensuring that our method can be easily deployed on a wide range of hardware platforms.
For network configuration, the Past Motion Encoder (PME), Future Motion Decoder (FMD), and Future Semantic Unit (FSU) each contain 3 stacked modules. The feature dimension is kept constant at 128 within PME and FMD, and at 256 within FSU. We employ 8-head attention, with each head containing 32 channels. During pretraining, the reconstruction losses of past and future motions share equal weights (=1). In the finetuning stage, the loss weight of text generation is set to 0.1, and the parameters of the FSU are updated every 5 epochs.
For self-supervised training on H36M, 3DPW, and AMASS, we train the model for 40 epochs and then finetune it for an additional 20 epochs on the motion prediction task. The best checkpoint on the validation set is selected after the self-supervised stage. For the processed FineMotion dataset, we train for 300 epochs in the self-supervised phase and subsequently finetune for 200 epochs on motion prediction.
9 More visualization results
We provide additional visualizations of motion prediction results on the processed FineMotion dataset in Fig. 7, and further show qualitative results on the Human3.6M dataset in Fig. 9 and Fig. 8. On Human3.6M, our method accurately predicts joint movements and maintains consistent body configurations over time. In particular, fine-grained actions involving limb articulations, such as arm swings, remain temporally coherent and physically plausible.
Compared with Human3.6M, the motions in FineMotion are more complex and the prediction horizon is significantly longer (up to 1.5 seconds). Even under these challenging conditions, our model successfully captures the correct motion tendencies and preserves the overall dynamics of large-amplitude arm movements and posture transitions. This robustness benefits from our self-supervised pretraining strategy, which encourages the model to learn implicit spatiotemporal constraints among joints. Nevertheless, we observe noticeable errors at distal joints, especially in highly articulated or fast-changing motion segments. This limitation partly arises from the inherent uncertainty and variability in the execution of fine-grained actions. In future work, we plan to address this issue by explicitly modeling motion uncertainty and incorporating stronger motion priors into the prediction framework.
@@ 10 Ablation Study on Model Architecture
Table 9: Ablation study on future text generation performance (future motion semantic understanding).
Method BLEU@4 BLEU@1 ROUGE CIDEr
w/o SSL 17.6 6.5 30.9 24.0
w/ SSL 18.2 7.8 32.5 24.2
Effect of SSL on Future Motion Understanding
We conduct an ablation study to examine how our self-supervised training stage influences the model’s ability to understand future motion semantics. As shown in Tab. 9, removing the SSL stage and directly training motion prediction and semantic understanding in a joint supervised manner (w/o SSL) leads to worse text generation performance. These results demonstrate that the SSL stage effectively enhances semantic reasoning about future dynamics and provides more informative motion representations for downstream captioning.
@@
Spatiotemporal Attention
In our default configuration, we adopt the sequential computation scheme from AuxFormer [xu2023auxiliary], where temporal attention and spatial attention are applied one after the other to progressively capture motion dynamics and joint dependencies. In contrast, the parallel computation variant (see Tab. 10 C), which processes temporal and spatial relations simultaneously, produces noticeably inferior performance, indicating that the sequential formulation provides a more effective inductive bias for structured human motion modeling.
Model Size
In Tab. 11, we show the impact of different model sizes (including feature dimension $d$ in attention layer and number of layers $l$) on parameter count and results. Although reducing the model size can decrease the number of parameters, it can also lead to decreased performance. We observe that the default setting (dd=256, ll=3) represents the optimal balance between parameter count and performance.
Figure 10: Visualization of global per-joint masking probability and the corresponding reconstruction error.
Figure 11: Visualization of reconstruction results. The blue lines are ground-truth value and the red lines represent the reconstruction motion skeleton sequences.
Table 12: Per-action reconstruction MPJPE at different time horizons ($m$).
Action 80ms 160ms 320ms 400ms
walking 2.55 6.25 12.54 15.48
eating 1.51 3.47 7.04 8.81
smoking 1.57 3.42 6.17 7.65
discussion 1.90 4.33 7.66 9.54
directions 1.45 3.41 7.39 9.45
greeting 2.45 6.08 12.52 15.64
phoning 1.88 4.47 9.28 11.63
posing 2.18 4.99 8.84 11.12
purchases 2.31 5.44 11.37 14.80
sitting 2.02 4.65 9.63 12.30
sittingdown 2.50 4.86 7.91 9.83
takingphoto 1.81 4.17 8.93 11.59
waiting 1.92 4.52 9.19 11.7
walkingdog 4.16 10.37 19.73 24.23
walkingtogether 1.91 4.89 10.92 14.00
Average 2.14 5.02 9.94 12.82
11 Analysis of Velocity-Based Joint Masking and Reconstruction Performance
To better understand how velocity-driven masking influences representation learning, we compute the global per-joint masking probability and the corresponding reconstruction error across the entire Human3.6M test set. The results are shown in Fig. 10. First, the masking distribution (left) exhibits a clear semantic pattern: High-velocity distal joints such as the feet, wrists, hands, and fingers are assigned higher masking probabilities; low-motion central joints, such as the spine, thorax, and neck, receive much smaller masking rates. This confirms that our velocity-based strategy automatically focuses learning pressure on dynamically informative regions of the body, rather than applying a uniform mask. Such non-uniform perturbation encourages the model to learn motion-consistent representations that reflect true physical dynamics, instead of memorizing static structure.
Second, the reconstruction error plot (right) shows a correlation with the masking distribution. Highly dynamic joints not only move faster but also demonstrate larger pose variability, making them inherently more challenging to reconstruct. Interestingly, although these dynamic joints are frequently masked, their reconstruction error does not explode. As shown in the right figure, the model maintains a stable reconstruction range (approximate 6 mm to 12 mm). This is largely because their adjacent, low-velocity joints remain visible and provide strong spatial constraints, allowing the model to leverage local kinematic continuity for recovery. In other words, the structured skeletal graph prevents error propagation and enables the model to infer masked high-frequency motions using nearby static or slow-moving joints.
The qualitative reconstructions in Fig. 11 further validate the effectiveness of our masking strategy. Even when large portions of the distal limbs are removed, the model is able to faithfully recover the full-body motion trajectory, producing temporally coherent poses that closely match the unmasked ground truth. Notably, complex motion cues—such as arm swings in Posing, subtle conversational gestures in Discussion, and periodic gait cycles in Walking—are accurately restored, demonstrating the model’s ability to infer fine-grained dynamics from limited visible context. These visual results highlight the robustness of our learned representations and showcase how motion structure priors guide reliable completion of challenging high-frequency movements.
The per-action results in Tab. 12 show that dynamic actions with large limb motions incur higher reconstruction errors, while static actions remain easy to recover. Despite this variation, our velocity-aware masking maintains stable performance across both low- and high-motion activities.